In a right triangle, the ratio of the two complementary angles is 5:4. What is the measure of the smaller of those two angles?
Answer by oberobic(2304)

The ratio of 5:4 means the 90 degrees of the complementary angles are divided into 9 parts.
90/9 = 10 degrees.
So, one angle is 5*10 = 50 degrees; and the other is 4*10 = 40 degrees.
50/40 = 5/4, which checks.
Answer: The smaller angle is 40 degrees.